# Detecting Edited Knowledge in Language Models

### v2

@@ Abstract
Knowledge editing methods (KEs) can update language models’ obsolete or inaccurate knowledge learned from pre-training. However, KEs can be used for malicious applications, e.g., inserting misinformation and toxic content. Knowing whether a generated output is based on edited knowledge or first-hand knowledge from pre-training can increase users’ trust in generative models and provide more transparency. Driven by this, we propose a novel task: detecting edited knowledge in language models. Given an edited model and a fact retrieved by a prompt from an edited model, the objective is to classify the knowledge as either unedited (based on the pre-training), or edited (based on subsequent editing). We instantiate the task with four KEs, two LLMs, and two datasets. Additionally, we propose using the hidden state representations and the probability distributions as features for the detection. Our results reveal that, using these features as inputs to a simple AdaBoost classifiers establishes a strong baseline. This classifier requires only a limited amount of data and maintains its performance even in cross-domain settings. Last, we find it more challenging to distinguish edited knowledge from unedited but related knowledge, highlighting the need for further research. Our work lays the groundwork for addressing malicious model editing, which is a critical challenge associated with the strong generative capabilities of LLMs.
@@ 1 Introduction
Large Language Models (LLMs) encode knowledge about the world via their pre-training data Petroni et al. (2019); Roberts et al. (2020). However, the encoded knowledge can be inherently flawed or become outdated with time De Cao et al. (2021). Additionally, practitioners might want to tailor LLMs by incorporating domain-specific knowledge pertaining to their products or to facilitate new applications. Driven by these needs and given the prohibitive cost of pre-training a new model, many methods have been developed to efficiently update knowledge in LLMs Meng et al. (2022); Zheng et al. (2023); Li et al. (2024a).
Knowledge editing methods (KEs), such as ROME (Meng et al., 2022), MEND (Mitchell et al., 2022a), and MEMIT (Meng et al., 2023), change the model’s internal weights to adapt facts. However, these methods can also be used to conduct malicious edits in LLMs Wang et al. (2023). For example, KEs can be used to inject fake news (as shown in Figure 1Halevy et al. (2024) or embed harmful or undesired biases in LLMs Mazzia et al. (2023). This danger becomes more pronounced as the number of capable open-source LLMs increases. While fact-checking and hate speech detection represent plausible avenues to combat these malicious uses (Zhao et al., 2021; Pan et al., 2023), these approaches require clearly defined objectives, as well as customized annotated datasets and models. Instead of utilizing specialized methods to address different issues, we take a vigilant approach: directly detecting edited knowledge. Moreover, detecting edited knowledge provides transparency to users in general application scenarios by revealing to them the source of LLM’s outputs (editd vs. unedited), and holds potential value for enhancing the controlled generation of LLMs (Pan et al., 2023; Sun et al., 2023).
Figure 1: Illustration of malicious knowledge editing, and our task of detecting edited knowledge (within the dashed box).
This work is the first to proposes the task of detecting edited knowledge in LLMs (DEED). We provide three detection baseline models and a comprehensive evaluation. We also consider cross-domain settings and challenging detection cases. Our experiments show that a simple AdaBoost classifier (Freund and Schapire, 1997) establishes a strong baseline when using hidden state representations and token probability distributions as input features. Our key contributions are as follows:
  • We propose a novel task for detecting edited knowledge (DEED) in LLMs to combat potential malicious model editing. DEED highlights the transparency of LLMs generations by suggesting the source of the generated text (edited vs. unedited) (Section 4).
@@
  • We show that simple AdaBoost classifiers, with the hidden state representations and probability distributions from the edited LLMs as features, establish strong baselines, while requiring a limited amount of training data (Section 5.2). Additionally, we show that these baselines generalize to cross-domain settings, highlighting their practical utility. Last, we find that distinguishing similar edited and unedited facts is more challenging, suggesting potential directions for future work (Section 6).
@@ 2 Related Work
This paper is the first to establish the task of detecting edited knowledge in pre-trained language models. Still, a wide array of work studies related questions, including how knowledge is stored in LLMs (Gurnee and Tegmark, 2023; Gurnee et al., 2024; Niu et al., 2024), how knowledge is probed (Youssef et al., 2023), how to inject new knowledge (Xu et al., 2023) and how to unlearn existing knowledge (Yu et al., 2023a; Kassem et al., 2023; Jang et al., 2023; Patil et al., 2024). For brevity, we direct readers to surveys by Mazzia et al. (2023) and Wang et al. (2023) for KE-related work above. This study focuses on the post-editing phase, i.e., after applying KEs that directly modify model parameters.
@@ 2.1 Knowledge Editing Methods
integrate extra modules, e.g., additional memory space or parameters, while leaving the original parameters unchanged (Yao et al., 2023; Li et al., 2024b). SERAC stores new knowledge in an explicit cache and uses an auxiliary classifier to determine if a prompt is associated with the edited knowledge. If it is, SERAC passes the prompt along with the stored new knowledge to a “counterfactual” model for the inference (Mitchell et al., 2022b). GRACE caches embeddings for old knowledge and the weights for new knowledge and adds an adaptor to manipulate the layer-to-layer transformations without altering model weights (Hartvigsen et al., 2023). MELO (Yu et al., 2023b) learns and stores new knowledge in inserted LoRA blocks (Hu et al., 2021). We exclude parameter-preserving KEs from our study, since for these KEs it is possible to undo the edits by simply removing the external modules from the model.
@@
modify the model parameters directly via meta-learning or optimization-based methods. Meta-learning KEs (ML-KEs) employ hyper-networks to learn parameter shifts towards new knowledge. KNOWLEDGE EDITOR (De Cao et al., 2021) and MEND (Mitchell et al., 2022a) build light-weight hyper-networks, i.e., naive multi-layer perceptrons that take the gradients of desired input-output (new knowledge) and output the parameter shifts for individual layers. However, the parameter shifts caused by different edits can be contradictory to each other, leading to a cancellation effect (Tan et al., 2024). To handle the cancellation effect in batch editing, MALMEN treats aggregating parameter shifts as a least square problem, seeking for the parameter shift to be effective for all edited facts (Tan et al., 2024).
On the other hand, optimization-based, also known as locate-and-edit methods (LE-KEs), assume that knowledge is memorized in a key-value form in the feed-forward network. These methods, such as ROME (Meng et al., 2022) and MEMIT (Meng et al., 2023), locate and then directly optimize the parameters in the feed-forward network to modify or add memories.
Note that fine-tuning the model on specific data roughly falls into the category of parameter-modifying KEs, but it requires great computational resources and data curation costs, and may encounter catastrophic forgetting and overfitting (Mitchell et al., 2022a, b; Zheng et al., 2023). These negative effects may result in an unusable model in practice. We leave detecting edits that have been induced through finetuning to future work, and focus on methods that have been specifically developed to edit knowledge.
@@ 2.2 Malicious Modifications of LLMs
With the development of KEs, knowledge of LLMs can be modified easily with low cost, and at scale. This development presents opportunities for misuse, such as inserting misinformation and toxic content into LLMs. For example, BadEdit by Li et al. (2024b) formulates backdoor injection as a lightweight knowledge editing problem, i.e., modifying a subset of parameters to inject backdoors into LLMs. Further, due to the high cost of training LLMs, users are likely to seek substitute models available on public resources such as GitHub and HuggingFace. Even though these third-party models might offer improved skills and advanced features (e.g., by fine tuning with Reinforcement Learning from Human Feedback (RLHF) or specific datasets), they could also be mingled with malicious edits before being made public Shi et al. (2023); Pan et al. (2023); Li et al. (2024b).
@@ 3 Preliminaries
Facts [^1] We use the terms facts and knowledge interchangeably. in LLMs are often represented as triplets of $(subject,relation,object)$, or $(s,r,o)$ for short. Querying an LLM with a prompt $p(s,r)$, where $p$ expresses the relation $r$ and contains the subject $s$ (e.g., “The Eiffel Tower is in the city of”), should result in retrieving the object $o$ (e.g., “Paris”), given that the fact $(s,r,o)$ is known to the LLM. A knowledge editing operation $E(s,r,o,o^{\prime},p)$ is successful if it changes the behavior of the LLM such that the retrieved object is $o^{\prime}$, as desired, instead of $o$. If an edit is conducted with a prompt $p$, then ideally it should also affect semantically similar (paraphrased) prompts that express that same relation, i.e., the new object $o^{\prime}$ should be retrievable with $p^{\prime}(s,r)$ as well, where $p^{\prime}$ is a paraphrase of $p$. We later use only paraphrased prompts to detect edited facts, since we assume no access to the prompts that have been used for editing.
@@ 4 DEED Task
The goal of detecting edited facts is to distinguish outputs generated based on post-hoc modifications to the model (edited knowledge) from those derived from the knowledge acquired by the model during its pre-training phase (unedited knowledge). In this section, we introduce the task of detecting edited knowledge in LLMs (DEED) in detail (Section 4.1), and describe how we instantiate the task (Section4.2).
@@ 4.1 DEED General Setup
More formally, we model the task of detecting edited knowledge as a binary classification problem with classes $\mathcal{Y}=\{edited,unedited\}$. Given an LLM $\mathcal{M}$ with unknown edits, and a pool of facts $\mathcal{T}_{test}=\{k_{1},..,k_{n}\}$, where $k_{i}=(s_{i},r_{i},o_{i},p_{i}),$ for $1\leq i\leq n$, the task is to assign each fact $k_{i}$ a label $\in\mathcal{Y}$.
In a supervised setting, the DEED classifier has access to a training set $\mathcal{T}_{train}$ with examples of edited and unedited facts. The facts are represented as a set of features $\mathcal{F}$. Conceptually, one can categorize the task based on the origin of the training set and the type of features; their combination defines the difficulty of the task.
Origin of the training set. If the training set is derived from the original model $\mathcal{M}$ (used also for testing), our task is In-Domain (ID) classification. When it originates from a similar LLM $\mathcal{M^{\prime}}$, which shares the same foundation model with $\mathcal{M}$, our task is Cross-Domain (CD) classification, e.g., $\mathcal{M}$ and $\mathcal{M^{\prime}}$ can be two different instruction-tuned versions of the same foundation model.
@@
A few state-of-the-art LLMs (e.g., GPT-4 Achiam et al. (2023)) are only accessible via an API. Nonetheless, the number of high-performing open-source LLMs is steadily increasing Liu et al. (2023b). The availability of these open-source LLMs allows users to further fine-tune and modify them locally, with access to their hidden states and probability distribution. In theory, the probability distribution can even be approximated for closed-source LLMs as well if enough outputs are sampled from these LLMs given a fixed input.
In the ID setting, the user has the training data for at least a subset of the edited facts made to $\mathcal{M}$, and these known edited facts are used to construct the training dataset. However, knowing some edited facts, while being unaware about other edited facts is highly unlikely in a real-world setting, which leads to our CD setting. In the CD setting, the user does not have such information about which facts have been edited. Nevertheless, the user can simulate the edit scenario by editing $\mathcal{M^{\prime}}$, an LLM of the same foundation model as $\mathcal{M}$, where the edited facts in $\mathcal{M^{\prime}}$ are known. Consequently, the edited facts in $\mathcal{M^{\prime}}$ are CD to the $\mathcal{T}_{test}$ extracted from $\mathcal{M}$.
@@ 4.2 DEED Constituents
We include four parameter-modifying KEs of two sub-categories, MEND (Mitchell et al., 2022a) and MALMEN (Tan et al., 2024) for meta-learning KEs, ROME (Meng et al., 2022) and MEMIT (Meng et al., 2023) for locate-and-edit KEs. We focus on parameter-modifying KEs, since for parameter-preserving KEs, it is possible undo the edits by simply removing the external modules from the model.
@@
We edit 1,000 facts in each setting (LLM, KE, dataset).222increasing the number of edits did not lead to more successful edits with all KEs. We consider 1,000 edits to be a realistic number for malicious editing attacks, and defer the investigation of this choice to future work. For ROME, we perform one edit at a time on the base language model as it is not designed for mass editing or sequential editing (Huang et al., 2022; Yao et al., 2023). For MEMIT, we edit each model with 1,000 facts at once. We use the hypernetworks for GPT-2 XL and GPT-J published by Meng et al. (2022) to circumvent the substantial compute required by MEND. For MEND, we also conduct one edit at a time to achieve a high success rate Meng et al. (2023). For MALMEN, we train the hyper-networks before editing and then perform 1,000 edits to the model simultaneously. We adhere to the default settings employed in the original papers. We include the unedited base model, NONE, as a yardstick. We follow related work Meng et al. (2022, 2023); Li et al. (2024a) in evaluating editing performance. See Appendix A for further details on editing metrics and performance.

### v3

# Has this Fact been Edited? Detecting Knowledge Edits in Language Models

@@ Abstract
Knowledge editing methods (KEs) can update language models’ obsolete or inaccurate knowledge learned from pre-training. However, KEs can be used for malicious applications, e.g., inserting misinformation and toxic content. Knowing whether a generated output is based on edited knowledge or first-hand knowledge from pre-training can increase users’ trust in generative models and provide more transparency. Driven by this, we propose a novel task: detecting knowledge edits in language models. Given an edited model and a fact retrieved by a prompt from an edited model, the objective is to classify the knowledge as either unedited (based on the pre-training), or edited (based on subsequent editing). We instantiate the task with four KEs, two large language models (LLMs), and two datasets. Additionally, we propose using hidden state representations and probability distributions as features for the detection model. Our results reveal that using these features as inputs to a simple AdaBoost classifier establishes a strong baseline. This baseline classifier requires a small amount of training data and maintains its performance even in cross-domain settings. Our work lays the groundwork for addressing potential malicious model editing, which is a critical challenge associated with the strong generative capabilities of LLMs. [^1] https://github.com/paulyoussef/deed
@@ 1 Introduction
Large Language Models (LLMs) encode knowledge about the world via their pre-training data Petroni et al. (2019); Roberts et al. (2020). However, the encoded knowledge can be inherently flawed or become outdated over time De Cao et al. (2021); Youssef et al. (2024). Additionally, practitioners might want to tailor LLMs by incorporating domain-specific knowledge pertaining to their products or to facilitate new applications. Driven by these needs and given the prohibitive cost of pre-training a new model, multiple methods have been developed to efficiently update knowledge in LLMs Meng et al. (2022, 2023); Li et al. (2024a).
Figure 1: Illustration of malicious knowledge editing, and our task of detecting knowledge edits. The cross-domain setting (dashed box) facilitates detection without known edits in the target LLM.
Knowledge editing methods (KEs), such as ROME (Meng et al., 2022), MEND (Mitchell et al., 2022a), and MEMIT (Meng et al., 2023), change the model’s internal weights to adapt facts. However, these methods can also be used to conduct malicious edits in LLMs Wang et al. (2023). For example, KEs can be used to inject fake news (cf. Figure  1) or embed harmful or undesired biases in LLMs Mazzia et al. (2023); Halevy et al. (2024); Youssef et al. (2025). This danger becomes more pronounced as the number of capable open-source LLMs increases. For example, any user can easily upload their model via one line of Python code to the HuggingFace Hub without validation and testing. While fact-checking and hate speech detection represent plausible avenues to combat these malicious uses (Zhao et al., 2021; Pan et al., 2023), these approaches require clearly defined objectives, as well as customized annotated datasets and models. Instead of utilizing multiple specialized methods to address the variety of malicious uses, we take a vigilant approach: directly detecting knowledge edits. Detecting knowledge edits additionally provides transparency to users in general application scenarios by revealing to them the source of LLM’s outputs (edited vs. unedited), and holds potential value for enhancing the controlled generation of LLMs (Pan et al., 2023; Sun et al., 2023). Moreover, detecting edits can help identify whether malicious outputs stem from poisoned pre-training data or from post-hoc artificial modifications.
This work is the first to propose the task of detecting knowledge edits in LLMs (DEED). We provide three detection baseline models and a comprehensive evaluation, including cross-domain settings and generalization across datasets. Our key contributions are as follows:
  • We propose a novel task for detecting knowledge edits in LLMs (DEED) to combat potential malicious model editing. DEED highlights the transparency of LLMs’ generations by suggesting the source of the generated text (edited vs. unedited) (Section 4).
@@
  • We show that simple AdaBoost classifiers, with the hidden state representations and probability distributions from the edited LLMs as features, establish strong baselines, while requiring a limited amount of training data (Section 5.2). Additionally, we show that these baselines generalize to cross-domain settings, highlighting their practical utility (Section 6).
@@ 2 Related Work
This paper is the first to establish the task of detecting knowledge edits in pre-trained language models. Still, a wide array of work studies related questions, including how knowledge is stored in LLMs (Gurnee and Tegmark, 2023; Gurnee et al., 2024; Niu et al., 2024), how knowledge is probed (Youssef et al., 2023), how to inject new knowledge (Xu et al., 2023) and how to unlearn existing knowledge (Yu et al., 2023a; Kassem et al., 2023; Jang et al., 2023; Patil et al., 2024). For brevity, we direct readers to surveys by Mazzia et al. (2023) and Wang et al. (2023) for KE-related work. This study focuses on the post-editing phase, i.e., after applying KEs that directly modify model parameters.
@@ 2.1 Knowledge Editing Methods
integrate extra modules, e.g., additional memory space or parameters, while leaving the original parameters unchanged (Yao et al., 2023; Li et al., 2024b). SERAC (Mitchell et al., 2022b) stores new knowledge in an explicit cache and uses an auxiliary classifier to determine if a prompt is associated with the edited knowledge. If it is, SERAC passes the prompt along with the stored new knowledge to a “counterfactual” model for the inference. GRACE (Hartvigsen et al., 2023) caches embeddings for old knowledge and the weights for new knowledge and adds an adaptor to manipulate the layer-to-layer transformations without altering model weights. MELO (Yu et al., 2023b) learns and stores new knowledge in inserted LoRA blocks (Hu et al., 2021). We exclude parameter-preserving KEs from our study, since for these KEs it is possible to undo the edits by simply removing the external modules from the model.
@@
modify the model parameters directly via meta-learning or optimization-based methods. Meta-learning KEs (ML-KEs) employ hyper-networks to learn parameter shifts towards new knowledge. KNOWLEDGE EDITOR (De Cao et al., 2021) and MEND (Mitchell et al., 2022a) build light-weight hyper-networks, i.e., naive multi-layer perceptrons that take the gradients of desired input-output (new knowledge) and output the parameter shifts for individual layers. However, the parameter shifts caused by different edits can be contradictory to each other, leading to a cancellation effect (Tan et al., 2024). To handle the cancellation effect in batch editing, MALMEN (Tan et al., 2024) treats aggregating parameter shifts as a least square problem, seeking for the parameter shift to be effective for all edited facts.
Optimization-based, also known as locate-and-edit methods (LE-KEs), assume that knowledge is memorized in a key-value form in the feed-forward network. These methods, such as ROME (Meng et al., 2022) and MEMIT (Meng et al., 2023), locate and then directly optimize the parameters in the feed-forward network to modify or add memories.
Note that fine-tuning the model on specific data roughly falls into the category of parameter-modifying KEs, but requires great computational resources and data curation costs, and may encounter catastrophic forgetting and overfitting (Mitchell et al., 2022a, b; Zheng et al., 2023). These negative effects may result in an unusable model in practice. We leave detecting edits that have been induced through finetuning to future work, and focus on methods that have been specifically developed to edit knowledge.
@@ 2.2 Malicious Modifications of LLMs
With the development of KEs, knowledge of LLMs can be modified easily with low cost, and at scale. This development presents opportunities for misuse, such as inserting misinformation and toxic content into LLMs. For example, BadEdit (Li et al., 2024b) formulates backdoor injection as a lightweight knowledge editing problem, i.e., modifying a subset of parameters to inject backdoors into LLMs. Due to the high cost of training LLMs, users are likely to seek readily available similar models on public resources such as GitHub and HuggingFace. Even though these third-party models might offer improved skills and advanced features (e.g., by fine tuning with Reinforcement Learning from Human Feedback (RLHF) or specific datasets), they could also be mingled with malicious edits before being made public Shi et al. (2023); Pan et al. (2023); Li et al. (2024b).
@@ 3 Preliminaries
Facts [^2] We use the terms facts and knowledge interchangeably. in LLMs are commonly represented as triplets of $(subject,relation,object)$, or $(s,r,o)$ for short. Querying an LLM with a prompt $p(s,r)$, where $p$ expresses the relation $r$ and contains the subject $s$ (e.g., “The Eiffel Tower is in the city of”), should result in retrieving the object $o$ (e.g., “Paris”), given that the fact $(s,r,o)$ is known to the LLM. A knowledge editing operation $E(s,r,o,o^{\prime},p)$ is successful if it changes the behavior of the LLM such that the retrieved object is $o^{\prime}$, as desired, instead of $o$. If an edit is conducted with a prompt $p$, then ideally it should also affect semantically similar (paraphrased) prompts that express that same relation, i.e., the new object $o^{\prime}$ should be retrievable with $p^{\prime}(s,r)$ as well, where $p^{\prime}$ is a paraphrase of $p$. We later use only paraphrased prompts to detect edited facts, since we assume no access to the prompts that have been used for editing. Paraphrased prompts are available in both datasets we consider.
@@ 4 DEED Task
The goal of detecting edited facts is to distinguish outputs generated based on post-hoc modifications to the model (edited knowledge) from those derived from the knowledge acquired by the model during its pre-training phase (unedited knowledge). In this section, we introduce the task of detecting knowledge edits in LLMs (DEED) in detail (Section 4.1), describe how we instantiate the task (Section 4.2), and discuss practical applications (Section 4.3).
@@ 4.1 DEED General Setup
More formally, we model the task of detecting knowledge edits as a binary classification problem with classes $\mathcal{Y}=\{edited,unedited\}$. Given an LLM $\mathcal{M}$ with unknown edits, and a pool of facts $\mathcal{T}_{test}=\{k_{1},..,k_{n}\}$, where $k_{i}=(s_{i},r_{i},o_{i},p_{i}),$ for $1\leq i\leq n$, the task is to assign each fact $k_{i}$ a label $y\in\mathcal{Y}$.
In a supervised setting, the DEED classifier has access to a training set $\mathcal{T}_{train}$ with examples of edited and unedited facts. The facts are represented as a set of features $\mathcal{F}$. Conceptually, one can categorize the task based on the origin of the test set and the type of features; their combination defines the difficulty of the task.
Origin of the test set. If the test set is derived from the original edited model, i.e., $\mathcal{M}=\mathcal{M}_{e}$ (used also for training), our task is In-Domain (ID) classification.When the test set originates from a similar LLM $\mathcal{M}=\mathcal{M}^{\prime}_{e}$, which shares the same foundation model with $\mathcal{M}_{e}$ (the model used to train the classifier), our task is Cross-Domain (CD) classification. For example, $\mathcal{M}_{e}$ and $\mathcal{M}^{\prime}_{e}$ can be two different instruction-tuned versions of the same foundation model (cf. Figure 1).
@@
A few state-of-the-art LLMs (e.g., GPT-4 Achiam et al. (2023)) are only accessible via their APIs. Nonetheless, the number of high-performing open-source LLMs is steadily increasing Liu et al. (2023b). The availability of these open-source LLMs allows users to further fine-tune and modify them locally, with access to their hidden states and probability distribution. In theory, the probability distribution can even be approximated for closed-source LLMs as well if enough outputs are sampled from these LLMs given a fixed input.
In the ID setting, the user knows at least a subset of the edited facts, and these known edited facts are used to construct the training dataset. However, in real-world settings we want to identify whether facts were edited and it is unrealistic to assume a known set of of edited facts. Therefore, we consider CD settings where the user does not know which facts have been edited in the model $\mathcal{M}=\mathcal{M}^{\prime}_{e}$. Nevertheless, users can simulate the edit scenario by editing $\mathcal{M}_{e}$, an LLM of the same foundation model as $\mathcal{M}^{\prime}_{e}$, where the edited facts in $\mathcal{M}_{e}$ are known. The Cross-domain box in Figure 1 illustrates this setting.
@@ 4.2 DEED Constituents
We include four parameter-modifying KEs of two sub-categories, MEND (Mitchell et al., 2022a) and MALMEN (Tan et al., 2024) for meta-learning KEs, and ROME (Meng et al., 2022) and MEMIT (Meng et al., 2023) for locate-and-edit KEs. We focus on parameter-modifying KEs, since for parameter-preserving KEs, it is possible to undo the edits by simply removing the external modules from the model.
@@
We edit 1,000 facts in each setting (LLM, KE, dataset). [^3] Increasing the number of edits did not lead to more successful edits with all KEs. We consider 1,000 edits to be a realistic number for malicious editing attacks, and defer the investigation of this choice to future work. For ROME, we perform one edit at a time on the base language model as ROME is not designed for mass editing or sequential editing (Huang et al., 2022; Yao et al., 2023). For MEMIT, we edit each model with 1,000 facts at once. We use the hypernetworks for GPT-2 XL and GPT-J published by Meng et al. (2022) to circumvent the substantial compute required by MEND. For MEND, we also conduct one edit at a time to achieve a high success rate Meng et al. (2023). For MALMEN, we train the hyper-networks before editing and then perform 1,000 edits to the model simultaneously. We adhere to the default settings employed in the original papers. We include the unedited base model, NONE, as a yardstick. We follow related work Meng et al. (2022, 2023); Li et al. (2024a) in evaluating editing performance. See Appendix A for further details on editing metrics and performance.
@@
4.3 Practical Applications
Detecting knowledge edits can have significant practical applications:
Transparency.
Detecting edits provides users with insights into whether model outputs originate from pre-training or post-deployment edits. This enhances the transparency of AI systems.
Model debugging.
By distinguishing between pre-training and edited knowledge, we can identify whether an issue stems from unclean pre-training data or a maliciously inserted edit. For example, if a factual error is flagged as an edit, developers can focus on investigating unauthorized modifications rather than pre-training data.
Mitigating malicious use cases.
Recent works highlight how knowledge editing can be used to bias LLMs Chen et al. (2024), introduce backdoors Li et al. (2024b); Qiu et al. (2024), or propagate misinformation in multi-agent settings Ju et al. (2024). Detecting knowledge edits helps identify and mitigate these threats by pinpointing unauthorized changes, ensuring the model’s reliability in adversarial environments.